Solve the system of three equations: a-b+6c=8, 3a-b+6c=14, -a-2b-3c=7. Write your answer as ordered triple.
Answer by checkley77(12844)

-[a-b+6c=8] subtract this equation from the one below.
3a-b+6c=14
----------------------
2a=6
a=6/2
a=3 ans.
-----
-a-2b-3c=7
-3-2b-3c=7 or -2b-3c=7+3 or -2b-3c=10
3*3-b+6c=14 or 9-b+6c=14 or -b+6c=14-9 or -b+6c=5 multiply by -2 & add.
2b-12c=-10
-2b-3c=10
-----------------
-15c=0
c=0 ans.
2b-3*0=-10
2b=-10
b=-10/2
b=-5 ans.